OpenStudy (anonymous):

25/100 =x/100 would that be 60% when the problem is solved?

OpenStudy (mathteacher1729):

Are you sure you copied the problem correctly? \[\frac{25}{100} = \frac{x}{100}\] The denominators are the same (they are both 100) so the numerators must be the same... that means x = 25.

OpenStudy (anonymous):

The orginal question is what percent of 25 is 15?

OpenStudy (anonymous):

25/15 x/100

OpenStudy (bahrom7893):

Set up a proportion: 25 - 100% 15 - x% x = (15*100)/25

OpenStudy (bahrom7893):

x = 60

OpenStudy (bahrom7893):

x = 60% I meant

OpenStudy (mathteacher1729):

Percentage is always "something over 100" 4/100 = 4 % 35/100 = 35 % and so on. So asking "what percentage of 25 is 15" is really asking: \[\frac{15}{25}=\frac{x}{100}\] From here you can cross multiply just like Bahrom did and have: \[x=\frac{15*100}{25}\] There is lots of canceling here, and you get x = 60, which means 60% in this case. :)